3. A man is 24_31/ older than his son. In two years, the man's age will be twice the age of his son. What will be the present age of his son?
Answer by macston(5194)

M=man's age ; S=son's age
M=S+(24_31/)
M+2=2(S+2) Substitute for M
S+(24_31/)+2=2(S+2)
S+(24_31/)+2=2S+2 Subtract S from each side.
(24_31/)+2=S+2 Subtract 2 from each side.
(24_31/)=S ANSWER: The son's age is (24_31/).